OpenStudy (anonymous):

In right ∆ABC, C is a right angle and sin A = 243/312 . What is m

jimthompson5910 (jim_thompson5910):

hint: find the angle A, then evaluate B = 90 - A

OpenStudy (anonymous):

how do you find angle A ?

jimthompson5910 (jim_thompson5910):

sin(A) = 243/312 A = arcsin(243/312) A = ???

OpenStudy (anonymous):

I got 89.99 but its not one of the answers

jimthompson5910 (jim_thompson5910):

not sine of 243/312 arcsine of that

jimthompson5910 (jim_thompson5910):

or inverse sine

jimthompson5910 (jim_thompson5910):

do you know what I mean?

OpenStudy (anonymous):

So wat the answers

OpenStudy (anonymous):

Because I got 89.99 idk that is not the answers

OpenStudy (anonymous):

Sin-1(243/312.)= 51.16 that us wgat @jim_thompdon5910 was telling you to do...